# Chevalley formulae for the motivic Chern classes of Schubert cells and for the stable envelopes

### v3

@@ Appendix B An example of the $\lambda$-chain formula
$\begin{array}[]{|c|c|c|c|c|c|}\hline\cr p_{w}\tilde{p}_{-\lambda}&\tilde{p}_{-\lambda}&\mathcal{M}&J&\varphi(p)=u&wt(p)=w\hat{r}_{J_{<}}(-\lambda)\\ \hline\cr\hline\cr(0)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{2}c^{+}_{1}c^{+}_{2}c^{+}_{0}c^{+}_{1}c^{+}_{2}}&\{\}&\{\}&s_{2}s_{1}&-\varpi_{1}+3\varpi_{2}\\ \hline\cr(1)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{2}c^{+}_{1}c^{+}_{2}c^{+}_{0}c^{+}_{1}}f^{-}_{2}&\{h_{6}\}&\{6\}&s_{1}&\varpi_{2}\\ (2)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{2}c^{+}_{1}c^{+}_{2}}f^{-}_{0}c^{+}_{1}c^{-}_{2}&\{h_{4}\}&\{4\}&s_{1}&\varpi_{1}-\varpi_{2}\\ (3)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{2}}f^{-}_{1}c^{+}_{2}c^{-}_{0}c^{+}_{1}c^{-}_{2}&\{h_{2}\}&\{2\}&s_{1}&2\varpi_{1}-3\varpi_{2}\\ \hline\cr(4)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{2}c^{+}_{1}c^{+}_{2}c^{+}_{0}}f^{-}_{1}c^{-}_{2}&\{h_{5}\}&\{5\}&s_{2}&-2\varpi_{1}+2\varpi_{2}\\ (5)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{2}c^{+}_{1}}f^{-}_{2}c^{-}_{0}c^{-}_{1}c^{-}_{2}&\{h_{3}\}&\{3\}&s_{2}&-3\varpi_{1}+\varpi_{2}\\ \hline\cr(6)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{2}}f^{-}_{1}f^{-}_{2}c^{-}_{0}c^{-}_{1}c^{-}_{2}&\{h_{2},h_{3}\}&\{2,3\}&id&-2\varpi_{1}-\varpi_{2}\\ (7)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{2}}f^{-}_{1}c^{+}_{2}c^{-}_{0}f^{-}_{1}c^{-}_{2}&\{h_{2},h_{5}\}&\{2,5\}&id&-2\varpi_{2}\\ (8)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{2}c^{+}_{1}c^{+}_{2}}f^{-}_{0}f^{-}_{1}c^{-}_{2}&\{h_{4},h_{5}\}&\{4,5\}&id&-\varpi_{1}\\ \hline\cr\end{array}$

### v4

@@ Appendix B An example of the $\lambda$-chain formula
$\begin{array}[]{|c|c|c|c|c|c|}\hline\cr p_{w}\tilde{p}_{-\lambda}&\tilde{p}_{-\lambda}&\mathcal{M}&J&\varphi(p)=u&wt(p)=w\hat{r}_{J_{<}}(-\lambda)\\ \hline\cr\hline\cr(0)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}c^{-}_{2}c^{+}_{1}c^{+}_{2}c^{+}_{0}c^{+}_{1}c^{+}_{2}}&\{\}&\{\}&s_{2}s_{1}&-\varpi_{1}+3\varpi_{2}\\ \hline\cr(1)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}c^{-}_{2}c^{+}_{1}c^{+}_{2}c^{+}_{0}c^{+}_{1}}f^{-}_{2}&\{h_{6}\}&\{6\}&s_{1}&\varpi_{2}\\ (2)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}c^{-}_{2}c^{+}_{1}c^{+}_{2}}f^{-}_{0}c^{+}_{1}c^{-}_{2}&\{h_{4}\}&\{4\}&s_{1}&\varpi_{1}-\varpi_{2}\\ (3)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}c^{-}_{2}}f^{-}_{1}c^{+}_{2}c^{-}_{0}c^{+}_{1}c^{-}_{2}&\{h_{2}\}&\{2\}&s_{1}&2\varpi_{1}-3\varpi_{2}\\ \hline\cr(4)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}c^{-}_{2}c^{+}_{1}c^{+}_{2}c^{+}_{0}}f^{-}_{1}c^{-}_{2}&\{h_{5}\}&\{5\}&s_{2}&-2\varpi_{1}+2\varpi_{2}\\ (5)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}c^{-}_{2}c^{+}_{1}}f^{-}_{2}c^{-}_{0}c^{-}_{1}c^{-}_{2}&\{h_{3}\}&\{3\}&s_{2}&-3\varpi_{1}+\varpi_{2}\\ \hline\cr(6)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}c^{-}_{2}}f^{-}_{1}f^{-}_{2}c^{-}_{0}c^{-}_{1}c^{-}_{2}&\{h_{2},h_{3}\}&\{2,3\}&id&-2\varpi_{1}-\varpi_{2}\\ (7)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}c^{-}_{2}}f^{-}_{1}c^{+}_{2}c^{-}_{0}f^{-}_{1}c^{-}_{2}&\{h_{2},h_{5}\}&\{2,5\}&id&-2\varpi_{2}\\ (8)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}c^{-}_{2}c^{+}_{1}c^{+}_{2}}f^{-}_{0}f^{-}_{1}c^{-}_{2}&\{h_{4},h_{5}\}&\{4,5\}&id&-\varpi_{1}\\ \hline\cr\end{array}$